# PET-MAD, a universal interatomic potential for advanced materials modeling

### v1

@@ Abstract
Machine-learning interatomic potentials (MLIPs) have greatly extended the reach of atomic-scale simulations, offering the accuracy of first-principles calculations at a fraction of the effort. Leveraging large quantum mechanical databases and expressive architectures, recent “universal” models deliver qualitative accuracy across the periodic table but are often biased toward low-energy configurations. We introduce PET-MAD, a generally applicable MLIP trained on a dataset combining stable inorganic and organic solids, systematically modified to enhance atomic diversity. Using a moderate but highly-consistent level of electronic-structure theory, we assess PET-MAD’s accuracy on established benchmarks and advanced simulations of six materials. PET-MAD rivals state-of-the-art MLIPs for inorganic solids, while also being reliable for molecules, organic materials, and surfaces. It is stable and fast, enabling, out-of-the-box, the near-quantitative study of thermal and quantum mechanical fluctuations, functional properties, and phase transitions. It can be efficiently fine-tuned to deliver full quantum mechanical accuracy with a minimal number of targeted calculations.
@@ I Introduction
Efficient methods for solving the electronic structure problem for molecules and materials have made it possible to predict their structure, stability and properties. Their computational cost and scaling with system size, however, put severe limits to the complexity, time, and length scale of the processes that can be simulated. Machine-learning (ML) models that aim to replace first-principles methods to assess energy and forces—so-called machine-learning interatomic potentials (MLIPs)—address this limitation by training on a small number of reference quantum mechanical calculations, and then making accurate yet inexpensive ML predictions, for both materials [1, 2, 3] and (bio)molecules [4, 5, 6, 7].
@@
For many years, this scheme has been successfully used to study a large number of chemical systems for which first-principles calculations are too expensive, and empirical force fields are either not available or not sufficiently accurate [8, 9, 10]. Nonetheless, the aforementioned approach requires the parametrization of a specific MLIP for every new system under study, which requires a considerable number of ab-initio calculations, as well as substantial human effort and expertise in fitting the potential. In contrast to this family of single-purpose potentials, recent years have seen the development of several general-purpose, or “universal” models [11, 12, 13] , which aim to be applicable, either out of the box or after minimal fine-tuning, to a large range of distinct chemical systems. These models provide a qualitative description of the atomic-scale interactions across the periodic table, although their level of accuracy for a specific application depends on how closely the problem of interest aligns with the considerations underlying the construction of the reference dataset. For instance, many existing datasets are built to represent a collection of stable materials, and to support the search for new materials with optimal properties.[14, 15, 16, 17, 18] To complicate things, the assessment of the accuracy of general-purpose potentials is often performed using inconsistent levels of ab-initio theory, making it difficult to isolate the shortcomings of the model from the discrepancy in the electronic-structure reference.
In this work, we introduce the Massive Atomistic Diversity (MAD) dataset, which incorporates a very high degree of chemical and structural diversity while employing highly converged, internally consistent first-principles calculations. This allows for a quantitative assessment of the performance of fitted models across the periodic table, since MAD’s ab-initio settings are broadly applicable to nearly any chemical system, although the comparatively simple exchange-correlation functional we use may not always be equally accurate. The energies and forces of the dataset are then fitted using the state-of-the-art Point Edge Transformer (PET) graph neural network [19], resulting in a general-purpose MLIP, which we name PET-MAD.
We evaluate the accuracy of the resulting model on a wide range of public datasets, recomputing the energies of a representative test set using consistent electronic-structure settings. We then critically assess the model’s performance in highly non-trivial atomistic simulations, including accelerated statistical sampling, quantum nuclear fluctuations, and the prediction of functional properties. We focus on six examples, motivated by their scientific interest, diversity in material classes and physical effects being probed, and the fact that each has previously been studied using ad hoc MLIPs—namely, lithium thiophosphate, gallium arsenide, a high-entropy alloy, liquid water, succinic acid and barium titanate. For each example, we assess the accuracy of PET-MAD by comparing it against both a bespoke model trained with compatible energetics, that provides a quantitative match to the electronic-structure reference, as well as a fine-tuned model that enables PET-MAD to achieve the same level of accuracy.
@@ II Results
This section introduces PET-MAD, a generally-applicable machine-learning interatomic potential based on PET (an unconstrained, transformer-based graph neural network architecture) and a custom training set built on the principles of Massive Atomistic Diversity (MAD) and internal consistency of the reference energetics. We thoroughly test the performance of PET-MAD against several benchmark datasets and compare it with four state-of-the-art universal potentials. Six case studies complete our analysis, showcasing and quantitatively assessing PET-MAD for diverse classes of materials and advanced atomistic simulations. We emphasize that in all of these examples we focus on the consistency of the model with the electronic-structure reference, irrespective of the actual accuracy of the latter in reproducing experimental quantities.
@@ II.1 MAD dataset
In contrast to previous universal models, which are either trained on datasets of inorganic crystals [14, 16, 17] or molecular datasets [20], PET-MAD is trained on a novel Massive Atomistic Diversity (MAD) dataset, which is based on a different philosophy. First, it aims to push the limits of “universality” by creating a model suitable for both organic and inorganic atomistic simulations in all possible system dimensionalities. Second, the resulting model is meant to work in complex atomistic simulation protocols describing a wide range of thermodynamic conditions, which requires covering a large configuration space. To this end, the MAD dataset includes randomized and nonequilibrium structures in the training set, applying to solids and surfaces ideas similar to those that drove the development of “mindless” molecular benchmark sets[21]. Third, the reference electronic-structure calculations are designed to be robust and consistent, to make sure that structural and chemical motifs are treated in the same way regardless of the type of structure they are part of, which is is important to ensure a coherent structure-energy mapping. This choice neglects the description of effects such as spin polarization, electron correlation, and dispersion, which are important for certain classes of materials but cannot be applied consistently across the MAD dataset. Details of the electronic-structure calculations are given in  Section IV.2. Last but not least, we restrict the dataset to fewer than 100,000 structures in order to reduce the time and cost of training, making the entire training procedure accessible to a wider community. As we shall see, the strategy we used allows us to preserve the representative power of the dataset.
@@
Surface slabs generated from the MC3D structures by cleaving the crystal along a random crystallographic plane with low-Miller index
@@ II.2 Model architecture and training
The PET architecture [19] is a rotationally unconstrained and transformer-based graph neural network (GNN) which offers excellent flexibility and scaling to large datasets. For PET-MAD, we chose the architecture based on an extensive hyperparameter search, resulting in a model with approximately 4 million parameters in total. Before training, each subset of MAD was randomly shuffled and further split into training, validation, and test subsets with fractions of 80%, 10%, and 10%, respectively. All subsets were then merged to obtain the final training, validation, and test sets. More details on the architecture and training are available in Section IV.3, Section IV.4 and the SI. In order to improve the performance of PET-MAD on specific chemical systems, we used a fine-tuning strategy based on the low-rank adaptation technique (LoRA) [27]. Details on fine-tuning are further discussed in Section IV.5.
@@
The model also includes uncertainty-quantification capabilities, based on the last-layer prediction rigidity (LLPR) method [28]. Error estimation is especially important for a model aiming to be generally-applicable, and that (despite the high diversity of the MAD training set) may often be used for out-of-domain predictions. As shown in Supplementary Figure S4, the calibrated LLPR uncertainties correlate well with the empirical test-set errors, while adding negligible overhead on top of energy predictions. Given that PET-MAD is intended for use in advanced simulations that estimate observable quantities through sophisticated sampling protocols, it is essential to be able to propagate the energy errors onto the final quantity of interest. To this end, we generate shallow ensembles based on the LLPR covariance. Unlike standard ensemble-based methods [29], which typically introduce a computational overhead proportional to the number of ensemble members, both LLPR and the shallow ensemble method [30] allow uncertainty quantification at no cost, while making it possible to use direct ensemble propagation to compute errors on any derived quantities, as we demonstrate in Section II.5. Details on our uncertainty-quantification approach are summarized in Section IV.6 and explained thoroughly in the Supplementary Text.
Figure 1: Comparison of accuracies of PET-MAD (orange line), MACE-MP-0-L (dark-blue line), Orb-v2 (light-blue line), MatterSim-5M (green line), and SevenNet-l3i5 (grey line) models on a subset of MAD dataset. The left and center panels show the mean absolute errors in energies and force predictions. The right panel shows the rotational discrepancy in energy predictions of the unconstrained models (PET-MAD and Orb-v2). Results for all models except PET-MAD dataset were obtained using a MAD subset recalculated with MPtrj dataset settings to achieve consistency with the training sets of the corresponding reference models.
@@ II.3 Benchmarking
We first assess the accuracy of the PET-MAD model on a subset of the MAD test set, comparing it to four recent universal MLIPs: MACE-MP-0 [11], MatterSim [12], Orb[13] and SevenNet [31]. Given that these models are trained relative to a different DFT setup (that of the MPtrj dataset [14]), and in order to perform a fair comparison, we restrict the chemical space of the MAD subset to the set of elements common to all models and recalculate energies and forces with the MPtrj settings (see the Supplementary Text for the selection details). We report the accuracy of each model relative to consistent reference values, evaluating each of the MAD subsets separately. We use the most advanced versions of the reference models: MACE-MP-0-L and MatterSim-5M, Orb-v2, SevenNet-l3i5, which have the largest number of parameters in their families. In evaluating this comparison, it is important to keep in mind that most of the models are larger than PET-MAD (having respectively about 15.8 M, 4.6 M, 25 M, and 1.17 M parameters), and are trained on much larger datasets (1.58 M, 17 M, 32.1 M, and 1.58 M structures, respectively). The results of this comparison are shown in Figure 1.
@@
Table 1: Comparison of PET-MAD accuracies on popular atomistic machine learning datasets against MACE-MP-0-L, MatterSim-5M, Orb-v2, and SevenNet-l3i5 models. For each dataset and model, mean absolute errors are reported for energy||||forces (in units of meV/atom||||meV/Å).
MAD MPtrj Alexandria SPICE MD22 OC2020
PET-MAD 19.4||||63.5 22.0||||79.1 52.2||||57.6 3.7||||59.8 1.9||||65.8 17.6||||114.6
MACE-MP-0-L 102.3||||182.8 16.0||||54.1 66.3||||70.5 10.6||||167.3 10.8||||219.5 82.8||||167.8
MatterSim-5M 60.6||||137.5 24.6||||64.3 25.7||||40.8 21.2||||146.2 25.8||||156.8 30.9||||118.2
Orb-v2 82.9||||101.4 5.5||||25.1 11.6||||9.1 70.8||||184.0 142.9||||321.7 20.6||||104.3
SevenNet-l3i5 108.2||||175.8 9.8||||29.0 40.8||||62.4 11.2||||141.2 13.4||||202.4 46.4||||157.5
Figure 2: Inference time of several models evaluated over different bulk materials and system sizes. For each MLIP, we use its LAMMPS [32] interface if available, preferably choosing the Kokkos-enabled [33] (kk) version, or its ASE [34] interface otherwise. All timings were measured on a single NVIDIA H100 GPU. Missing points for MACE-MP-0 and SevenNet caused by out-of-memory errors are marked with red crosses. We note that Orb is a non-conservative model, and therefore benefits from a theoretical speed-up; however, the violation of conservation of energy often results in ill-behaved molecular dynamics [35]. The model versions used were MatterSim-v1.0.0-1M, MACE-MP-0 (M), Orb-v2, SevenNet-0 (11Jul2024).
PET-MAD outperforms the MACE-MP-0-L, MatterSim-5M and SevenNet-l3i5 models in almost all cases. On bulk inorganic crystal subsets (MC3D, MC3D-rattled and MC3D-randcomp), PET-MAD and MatterSim-5M show similar accuracy, while MACE-MP-0-L and SevenNet-l3i5 give significantly worse results on rattled and random composition structures. These observations are not surprising, since the MPtrj dataset, which MACE-MP-0-L and SevenNet-l3i5 is trained on, consists mainly of relaxation trajectories of ideal bulk crystals and thus covers a smaller fraction of the configuration space than MAD. On the contrary, MatterSim-5M is trained on a broader, yet unpublished, set of configurations, which presumably leads to a better coverage of the configuration space. On the subsets of surface slabs (MC3D-surface), 2D crystals (MC2D), and nanoclusters (MC3D-cluster), the models show similar performance patterns, but this time PET-MAD manages to better capture the energies of the low-dimensional structures, which are explicitly included in the training set, except for the case of SevenNet that performs slighly better for the MC2D subset. The most notable difference occurs in the molecular systems (SHIFTML-molcrys and SHIFTML-molfrags), where PET-MAD dramatically outperforms MACE-MP-0-L and MatterSim. This result is again expected, as organic systems represent a completely different region of configuration space, which is heavily undersampled in the case of inorganic crystal datasets.
The last model from the reference list, Orb-v2, requires a separate discussion, since it represents a class of so-called non-conservative models, which are becoming increasingly popular. Compared to conservative PET-MAD, MACE-MP-0, MatterSim, and SevenNet, these models directly predict the energies and forces of the system as two independent properties, so that the forces are no longer derivatives of the energies w.r.t. the positions of the system. As a consequence, these models usually have faster inference and training, but they lead to serious artifacts when used in molecular-dynamics simulations[35]. Orb-v2 outperforms all other models (including, marginally, PET-MAD) on relaxed inorganic systems, presumably thanks to the combination of a flexible, unconstrained architecture and the large training set (which includes both MPtrj [14] and Alexandria [16, 17] datasets). However, it still provides worse accuracy on rattled, random composition structures, surfaces and molecular systems compared to PET-MAD. Both PET-MAD and Orb-v2 use architectures that do not enforce exact rotational equivariance, which is learned approximately by data augmentation. Even though the resulting symmetry breaking can be monitored and controlled easily [36], it is important to assess its extent. To do so, we estimated the symmetry error in PET-MAD and Orb-v2 energy predictions by applying a series of rotations on a Lebedev-Laikov grid of order 9 for each structure in the MAD subset and calculating the standard deviation in the predictions. The corresponding results are presented in the right panel of Figure 1. In most cases, the rotational discrepancy of the PET-MAD model is one or two orders of magnitude smaller than the corresponding prediction error, below 1 meV/atom for all subsets except MC3D-random and MC3D-cluster. Orb-v2 shows significantly higher discrepancies in energy predictions, which are at times comparable to the actual errors.
The superior performance of PET-MAD on its own test set is not entirely surprising – even though one should consider the much smaller size of the MAD dataset in comparison to the training sets used for the other models. As a more challenging test, we evaluate the accuracy of PET-MAD on some of the most popular atomistic machine learning benchmarks, which include bulk inorganic systems, molecular systems, and catalytic applications, i.e.:
@@
Relaxation trajectories of bulk inorganic crystals dataset from Ref. 14
@@
Relaxation trajectories of bulk inorganic crystals as well as 2D, 1D systems from Ref. 16 and 17
@@
Drug-like molecules and peptides from Ref. 20
@@
OC2020 (S2EF)
Molecular relaxation trajectories on catalytically active surfaces from Ref. 38
In order to maintain consistency between the PET-MAD training set DFT settings and the target benchmark data, we recalculated a small fraction of each dataset from the list above using our DFT settings and used them to evaluate the accuracy of PET-MAD (the details on subset selection are provided in the Supplementary Text). We also include the results of the evaluation of the MACE-MP-0-L, MatterSim-5M, Orb-v2 and SevenNet-l3i5 models for comparison. Here we follow the same rule of consistency and recalculate the molecular benchmarks (SPICE and MD22) and the OC2020 subset using the MPtrj dataset DFT settings, while keeping the original data on energies and forces for the MPtrj and Alexandria subsets, since they are already consistent.

### v2

# PET-MAD, a lightweight universal interatomic potential for advanced materials modeling

@@ Abstract
Machine-learning interatomic potentials (MLIPs) have greatly extended the reach of atomic-scale simulations, offering the accuracy of first-principles calculations at a fraction of the cost. Leveraging large quantum mechanical databases and expressive architectures, recent “universal” models deliver qualitative accuracy across the periodic table but are often biased toward low-energy configurations. We introduce PET-MAD, a generally applicable MLIP trained on a dataset combining stable inorganic and organic solids, systematically modified to enhance atomic diversity. Using a moderate but highly-consistent level of electronic-structure theory, we assess PET-MAD’s accuracy on established benchmarks and advanced simulations of six materials. Despite the small training set and lightweight architecture, PET-MAD is competitive with state-of-the-art MLIPs for inorganic solids, while also being reliable for molecules, organic materials, and surfaces. It is stable and fast, enabling the near-quantitative study of thermal and quantum mechanical fluctuations, functional properties, and phase transitions out of the box. It can be efficiently fine-tuned to deliver full quantum mechanical accuracy with a minimal number of targeted calculations.
@@ I Introduction
Efficient methods for solving the electronic structure problem for molecules and materials have made it possible to predict their structure, stability and properties from first principles. However, the associated computational cost and poor scaling with system size severely limit the complexity, time, and length scale accessible to simulations. Machine-learning (ML) models that aim to replace first-principles methods to evaluate energy and forces — so-called machine-learning interatomic potentials (MLIPs) — address this limitation by training on a set of reference quantum mechanical calculations, and then making accurate yet inexpensive ML predictions, for both materials [2, 3, 4] and (bio)molecules [5, 6, 7, 8].
@@
For almost two decades, this scheme has been used successfully to study a large number of chemical systems for which first-principles calculations are too expensive, and empirical forcefields are either not available or not sufficiently accurate [9, 10, 11]. Nonetheless, the aforementioned approach typically implies parameterizing a specific MLIP for every new system, which requires a considerable number of ab initio calculations, as well as substantial human effort and expertise in fitting the potential. In contrast to this family of single-purpose potentials, recent years have seen the development of several general-purpose, or “universal” models [12, 13, 14], which aim to be applicable, either out of the box or after minimal fine-tuning, to a large range of distinct chemical systems. These models provide a qualitative description of the atomic-scale interactions across the periodic table, although their level of accuracy for a specific application depends on how closely the problem of interest aligns with the considerations underlying the construction of the reference dataset. For instance, many existing datasets are built to represent a collection of stable materials, and to support the search for new materials with optimal properties [15, 16, 17, 18]. To complicate things, the accuracy of general-purpose potentials is often assessed using inconsistent levels of ab initio theory, making it difficult to isolate the shortcomings of the model from the discrepancy in the electronic-structure reference. In this work, we use the Massive Atomistic Diversity (MAD) dataset, which incorporates a high degree of chemical and structural diversity while employing highly converged, internally consistent reference calculations. This allows for a quantitative assessment of the performance of fitted models across the periodic table, since MAD’s ab initio settings are broadly applicable to most chemical systems, although the comparatively simple exchange-correlation functional we use may not always be equally accurate. Energies and forces are fitted using a Point Edge Transformer (PET) graph neural network [19], resulting in a general-purpose MLIP, which we name PET-MAD. We evaluate the accuracy of the resulting model on a wide range of public datasets, ensuring consistency in the electronic-structure settings, and showing that PET-MAD achieves competitive performance on several external benchmarks despite being trained on 2-3 orders of magnitude fewer structures. We then critically assess the performance of PET-MAD in highly non-trivial atomistic simulations, including accelerated statistical sampling, quantum nuclear fluctuations, and predicting functional properties. We focus on six examples, motivated by their scientific interest, diversity in material classes and physical effects being probed, and the fact that each has previously been studied using ad hoc MLIPs — namely, lithium thiophosphate, gallium arsenide, a CoCrFeMnNi high-entropy alloy, liquid water, succinic acid, and barium titanate. For each example, we quantitatively assess the accuracy of PET-MAD by comparing it against both a bespoke model trained with compatible energetics, that provides an extremely close match to the electronic-structure reference, as well as a finetuned model that enables PET-MAD to achieve the same level of accuracy with a small number of additional training structures. PET-MAD demonstrates that accurate, fast, and robust universal models can be trained using a tiny fraction of the structures of last-generation datasets, and provides a feature-rich framework for advanced atomistic simulations, which includes direct-force acceleration and inexpensive end-to-end uncertainty quantification.
@@ II Results
PET-MAD is a generally-applicable machine-learning interatomic potential based on PET (an unconstrained, transformer-based graph neural network architecture) and a custom training set built on the principles of Massive Atomistic Diversity (MAD) and internal consistency of the reference energetics. We thoroughly test the performance of PET-MAD against several benchmark datasets and compare it with four widely used universal potentials. Six case studies complete our analysis, showcasing and quantitatively assessing PET-MAD for diverse classes of materials and advanced atomistic simulations, comparing a system-specific PET versions trained on an ad hoc datasets, the generally-applicable PET-MAD, and the versions finetuned to each system. We emphasize that in all of these examples we focus on the consistency of the model with the electronic-structure reference, irrespective of the actual accuracy of the latter in reproducing experimental quantities.
@@ II.1 MAD dataset
Most of the existing efforts to generate datasets to train universal models focus on either inorganic crystals [15, 20, 17] or molecular compounds [21], and aim to include as many structures as possible. PET-MAD is trained on the Massive Atomistic Diversity (MAD) dataset, which is based on a different philosophy. First, it aims to push the limits of “universality” by including both organic and inorganic atomistic systems of all possible dimensionalities. This approach has been already shown to deliver good transferability upon including up to 45 elements in the training data [22]. Second, our resulting model is meant to work in complex atomistic simulation protocols describing a wide range of thermodynamic conditions, which requires covering a large configuration space and being computationally affordable. To this end, the MAD dataset includes randomized and highly-distorted structures in the training set, applying to solids and surfaces ideas similar to those that drove the development of “mindless” molecular benchmark sets [23]. Third, the reference electronic-structure calculations are designed to be robust and consistent, to ensure that the structural and chemical motifs are treated in the same way, regardless of the type of structure they are part of, which is important for a coherent structure-energy mapping. This choice neglects the description of effects such as spin polarization, electron correlation, and dispersion, which are important for certain classes of materials but cannot be applied consistently across the MAD dataset. Details of the electronic-structure calculations are given in Section IV.2. Last but not least, we restrict the dataset to fewer than 100,000 structures in order to reduce the time and cost of training, making the entire training procedure accessible to a wider community. As we shall see, the strategy we use allows us to preserve the representative power of the dataset.
@@
Surface slabs generated from the MC3D structures by cleaving the crystal along a random crystallographic plane with low Miller index (5589 structures)
@@ II.2 Model architecture and training
The PET architecture [19] is a rotationally unconstrained and transformer-based graph neural network (GNN) which has a high descriptive power (as every transformer layer can be scaled to be a universal approximator) [19] and low inference cost (as we shall demonstrate in Fig. 3). For PET-MAD, we chose the architecture based on an extensive hyperparameter search, resulting in a model with approximately 3.3 million parameters in total. Before training, each subset of MAD was randomly shuffled and further split into training, validation, and test subsets in fractions of 80%, 10%, and 10%, respectively. All subsets were then merged to obtain the final training, validation, and test sets. More details on the architecture and training are available in Section IV.3, Section IV.4, and the Supplementary Information. In order to improve the performance of PET-MAD on specific chemical systems, we used a fine-tuning strategy based on the low-rank adaptation technique (LoRA) [29]. Details on fine-tuning are further discussed in Section IV.5.
@@
The model also includes uncertainty-quantification capabilities, based on the last-layer prediction rigidity (LLPR) method [30]. Error estimation is especially important for a model aiming to be generally-applicable, and that (despite the high diversity of the MAD training set) may often be used for out-of-domain predictions. As shown in Supplementary Figure S6, the calibrated LLPR uncertainties correlate well with the empirical test-set errors, while adding negligible overhead on top of energy predictions. Given that PET-MAD is intended for use in advanced simulations that estimate observable quantities through sophisticated sampling protocols, it is essential to be able to propagate the energy errors onto the final quantity of interest. To this end, we generate a shallow ensemble based on the LLPR covariance [30]. Unlike standard ensemble-based methods [31], which typically introduce a computational overhead proportional to the number of ensemble members, both LLPR and the shallow ensemble method [32] allow uncertainty quantification at negligible additional cost. Moreover, they allow one to compute errors on any derived quantities using direct ensemble propagation, as we demonstrate for free-energy calculations in Section II.5, and for phonon dispersion curves in Section S8. Details on our uncertainty-quantification approach are summarized in Section IV.6 and explained thoroughly in the Supplementary Information. Even though PET-MAD is designed to evaluate forces using backpropagation, we also train a separate head that predicts forces directly as a function of the atomic coordinates. Avoiding backpropagation makes the model two to three times faster (Figure 3), but should be used with care, as it violates energy conservation with pathological consequences on sampling accuracy [33]. We address these problems in Section S12, where we also show how a multiple-time-step integrator allows one to retain the computational advantages of direct force predictions without introducing sampling artifacts.
Figure 1: Pareto frontier of various universal MLIPs: x-axis shows the number of configurations in the training set, and y-axis shows mean absolute error (MAE) for the energy-above-hull property on a subset of the WBM dataset [16] (a part of the Matbench Discovery benchmark [34]). Colored dots represent a set of selected models, while other models from the benchmark are shown with grey dots. PET-MAD results are shown with diamonds for both consistent and inconsistent DFT references. PET-MAD significantly pushes the frontier downward, achieving the accuracy of other advanced universal MLIPs like SevenNet and GNoME using 1-3 orders of magnitude less data.
@@ II.3 Benchmarking
First, we demonstrate the efficacy of combining the MAD dataset and the PET architecture by comparing the accuracy of PET-MAD against other universal models on the popular Matbench Discovery benchmark [34]. We analyze the overall gain in models’ accuracies with the increase in the training set size and the number of trainable parameters.
Since PET-MAD is trained on the MAD dataset, which has different DFT settings compared to the Matbench (which shares common settings with the training sets of other reference models), we recomputed a random subset of the benchmark containing 555 structures (excluding lanthanides and actinides) using MAD-compatible DFT settings. We evaluate the accuracies of all models on this subset, picking for each model the DFT reference compatible with its training set. Figure 1 demonstrates the Pareto frontier of the models’ data efficiency: Pareto-optimal models efficiently use additional training data to increase their accuracy. PET-MAD significantly improves on existing models, achieving the accuracy of other advanced universal MLIPs like SevenNet and GNoME using 1-3 orders of magnitude less data. For illustrative purposes, the figure also shows the accuracy of PET-MAD evaluated against the inconsistent Matbench energetics: the $3\times$ larger error is entirely due to the discrepancy in the reference data. A thorough analysis of DFT consistency effects is discussed in the Supplementary Section S3
@@
Table 1: Comparison of PET-MAD accuracies on popular atomistic machine learning datasets against MACE-MP-0-L, MatterSim-5M, Orb-v2, and SevenNet-l3i5 models. For each dataset and model, mean absolute errors are reported for raw energy||forces predictions (in units of meV/atom||meV/Å). We do not show force errors for the Matbench Discovery data, since they are not available in the reference data.
Dataset PET-MAD MACE-MP-0-L MatterSim-5M Orb-v2 SevenNet
MAD 17.6||65.1 81.6||181.5 47.3||133.7 52.9||96.2 82.1||173.5
MPtrj 22.3||77.9 15.1||50.8 21.3||61.4 5.6||21.9 9.8||25.5
Matbench 31.3|| 58.5|| 38.2 $|$ 37.9|| 47.5||
Alexandria 49.0||66.8 65.4||79.5 21.2||39.9 13.2||10.5 47.6||70.3
OC2020 18.3||114.5 82.4||169.6 31.5||119.2 19.8||99.3 45.7||162.7
SPICE 3.7||59.5 10.6||166.8 21.3||145.6 59.0||140.8 11.3||139.1
MD22 1.9||65.6 9.4||182.9 28.6||160.4 174.3||220.7 11.1||146.2
The figure also shows that many of the optimal models use a very large number of parameters, which makes them accurate but computationally demanding, and therefore not suitable for the long, complex simulation tasks that PET-MAD is designed for. For this reason, we will focus on medium-size models for the rest of these benchmarks: MACE-MP-0 [12], MatterSim [13], Orb-v2 [14] and SevenNet [35]. Like PET-MAD, these models have been designed and are widely used for advanced materials simulations beyond single-point calculations. We extend the comparison on Matbench to include other popular atomistic machine learning benchmarks for bulk inorganic systems, molecular systems, and catalytic applications. Thus, we use a total of seven following benchmarks:
@@
Relaxation trajectories of bulk inorganic crystals dataset from Ref. 15
Matbench
The dataset of single-element substitutions on bulk inorganic crystals from Ref. 16
@@
Relaxation trajectories of bulk inorganic crystals as well as 2D, 1D systems from Ref. 20
OC2020 (S2EF)
Molecular relaxation trajectories on catalytically active surfaces from Ref. 36
@@
Drug-like molecules and peptides from Ref. 21
@@
In evaluating the benchmarks, it is important to keep in mind that (1) most of the reference models are larger than PET-MAD (we use the versions with the largest number of parameters in their families, MACE-MP-0-L and MatterSim-5M, Orb-v2, SevenNet-l3i5 having respectively about 15.8 M, 4.6 M, 25 M, and 1.17 M parameters, as opposed to 3.3 M for PET-MAD) and (2) they are trained on much larger datasets (1.58 M, 17 M, 32.1 M, and 1.58 M structures, respectively).